I need to evaluate the following expressions:
log of base 5 (0.008)

and
log of base 64 sqrt(4)

Answer by stanbon(48535)

I need to evaluate the following expressions:
log of base 5 (0.008)
=ln(0.008)/ln5
=-4.828313737.../1.609437912...= -3
----------------
and
log of base 64 sqrt(4)
=ln(sqrt(4))/ln(64)
=0.693147.../4.158883...
= 0.1666666666667
---------
Cheers,
Stan H.